What is the electrostatic force between and electron and a proton separated by 0.1 mm?
Answer #1

Here ,

for the electrostatic force

electrostatic force = k * e^2/d^2

electrostatic force = 9 *10^9 * (1.602 *10^-19)^2/(1 *10^-4)^2

electrostatic force = 2.31 *10^-20 N

the electrostatic force between them is 2.31 *10-20 N